1. (DCH Page 23) Cupric sulfate is commonly sold as the pentahydrate, CuSO4-H2O, which has 5 moles of H20 for each mole of CuSO4 in the solid. The formula mass is 249.69 g/mol. How many grams of CuS04 H2O should be dissolved in a 250 mL volumetric flask to make a solution containing 8.00 mM Cu2* (a) 1.28 g (b) 0.499 g (c) 2.00 g (d) 0.319 g (e) 3.72 g
Answer #1

Molarity = (W /MW) (1000 / V in mL)

Molarity = 8.00 mM = 0.008 M

0.008 = (W / 249.69) (1000 / 250)

W / 249.69 = 0.002

W = 0.499 g

answer = option b = 0.499 g